# Streaming Neural Images

### v1

@@ 3.2 Comparison with traditional Codecs
Fig. 4: Comparison of INRs with standard codecs [7, 17]. The baseline INRs represent a constant bpp level.
@@ 3.3 Model Complexity
In Table 1 we study models with $h=256,l=2$ (highlighted with “*”) and smaller models with $h=128,l=4$ (ie., half width, double depth). We can conclude that only MLP-based approaches with “low” complexity can be good image compressors, offering a positive compression factor (CF). For instance CF$>2$ indicates the INR is twice smaller than the natural image. See these models also in Fig. 4.
@@ 3.4 Unstable Training and Sensitivity
FourierNet* [1] 132.09 30.78$\pm$0.23 0.857 1.22
SIREN* [2] 133.12 31.68$\pm$2.08 0.866 1.22
MFN* [9] 136.96 33.98$\pm$0.18 0.909 1.17
FourierNet [1] 66.30 30.10$\pm$0.42 0.832 2.41
SIREN [2] 66.82 30.95$\pm$2.31 0.855 2.40
MFN [9] 70.27 32.95$\pm$0.47 0.895 2.25
Wire [10] 66.82 25.96$\pm$3.83 0.712 1.20
DINER [11] 545.55 27.34$\pm$15.5 0.751 0.32
@@ 4 SPINR: Streamable Progressive INR
0:  Input $x$, Ground Truth $y$, Stages $S=[1,n+1]$, Number of optimization steps $N$
@@
Considering $n+1$ stages, in stage 1 we train the smallest possible mapping: $\mathbf{y}=L_{5}\circ L_{0}(\mathbf{x})$. In the next stage, we freeze the previously trained layers (eg., [$L_{0},L_{5}$]) and train the next connection $\mathbf{y}=L_{5}\circ L_{1}\circ L_{0}(\mathbf{x})$, updating only $L_{1}$ from the loss backpropagation. We repeat this process until we train $L_{n}$ having as frozen layers [$L_{0},...,L_{(}n-1),L_{5}$]. This allows the model to learn the $\mathbf{x}\to\mathbf{y}$ mapping considering different connections (forward pass) within the INR.

### v2

@@ 3.2 Comparison with traditional Codecs
00.10.10.20.20.30.30.40.40.50.50.60.60.70.70.80.80.90.9112020252530303535Bitrate [bpp]PSNR [dB]JPEGJPEG2000COIN [3]COIN++ (5-bit) [5]SIREN (8-bit) [4]
Fig. 4: Comparison between INR compression [4, 3, 5] and classical JPEG compression using the Kodak dataset. Best viewed in electronic version.
@@ 3.3 Model Complexity
In Table 1 we study models with $h=256,l=2$ (highlighted with “*”) and smaller models with $h=128,l=4$ (ie., half width, double depth). We can conclude that only MLP-based approaches with “low” complexity can be good image compressors, offering a positive compression factor (CF). For instance CF>2>2 indicates the INR is twice smaller than the natural image. See these models also in Fig. 4.
@@ 3.4 Unstable Training and Sensitivity
[…]
@@ 4 SPINR: Streamable Progressive INR
0: Input $x$, Ground Truth $y$, Stages $S=[1,n+1]$, Number of optimization steps NN
@@
Considering $n+1$ stages, in stage 1 we train the smallest possible mapping: $\mathbf{y}=L_{5}\circ L_{0}(\mathbf{x})$. In the next stage, we freeze the previously trained layers (eg., $L_{0},L_{5}$]) and train the next connection $\mathbf{y}=L_{5}\circ L_{1}\circ L_{0}(\mathbf{x})$, updating only $L_{1}$ from the loss backpropagation. We repeat this process until we train $L_{n}$ having as frozen layers $L_{0},...,L_{(}n-1),L_{5}$]. This allows the model to learn the $\mathbf{x}\to\mathbf{y}$ mapping considering different connections (forward pass) within the INR.